OpenStudy (anonymous):

The table shows the solution to the equation |2x - 5| - 2 = 3: Step 1 |2x - 5| = 3 + 2 Step 2 |2x - 5| = 5 Step 3 2x - 5 = 5 or 2x - 5 = -5 Step 4 2x = 10 or 2x = 0 Step 5 x = 5 or x = 0 Which is the first incorrect step? Step 3 Step 4 Step 5 Solution is correct

OpenStudy (anonymous):

Try solving this Q yourself. I'll help you out wherever you need it.

OpenStudy (anonymous):

I don't think any are.

OpenStudy (anonymous):

So you think it's D?

OpenStudy (anonymous):

I think so.

OpenStudy (anonymous):

Let's see… (Give me 2 mins to solve it and show you)

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

Clear the absolute-value bars by splitting the equation into its two cases, one for the Positive case and the other for the Negative case. The Absolute Value term is |2x-5| For the Negative case we'll use -(2x-5) For the Positive case we'll use (2x-5)

OpenStudy (anonymous):

Is everything clear so far?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Solve the Negative Case -(2x-5) = 5 Multiply -2x+5 = 5 Rearrange and Add up -2x = 0 Divide both sides by 2 -x = 0 Multiply both sides by (-1) x = 0 Which is the solution for the Negative Case

OpenStudy (anonymous):

Distribution is aka Multiply

OpenStudy (anonymous):

Solve the Positive Case (2x-5) = 5 Rearrange and Add up 2x = 10 Divide both sides by 2 x = 5 Which is the solution for the Positive Case

OpenStudy (anonymous):

Clear?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Wrap up the solution x=0 x=5

OpenStudy (anonymous):

So, your initial answer is correct! Anything you'd like me to explain more for?

OpenStudy (anonymous):

Nope, u made it crystal clear.

OpenStudy (anonymous):

Awesome to know! :)

OpenStudy (mathstudent55):

Just follow the steps in the problem. Every step is correct, so there are no mistakes. I don't see the reason to complicate this problem.